An Australian firm has business relation with some European companies in Euro zone and Thailand, so CFO is concerned about Euro and Thai Baht (THB) value relative to Australian dollar. The firm is given the following quote: AUD/USD 0.7980 EUR/USD 1.1990 USD/THB 32.50 What is the exchange rate for EUR/AUD and AUD/THB? EUR/AUD EUR/USDAUD/USD 1.19900.7980 1.5025 AUD/THB AUD/USD x USD/THB 0.7980 x 32.50- 25.9350

Why divide one and multiply another?
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text ta copy. в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 Conditional Format eCell Insert Delete Format Paste Sort &Find & 2 ClearFe Select Edting Format Painter Formatting, as Table w styles. Styles ▼ ㆆ ▼ Clipboard Font Alignment Number Cells BA59 AP AQ AR AS AT AU AV AW AX AY AZ BA BC BD BE 59 60 61 62 63 64 65 IT IS A SUM OF CROSS RATES IT INVOLVES BASIC MATHS FIRST ONE EUR/AUD WE HAVE THREE CURRENCIES EUR, AUD, THB AS EUR/AUD IS NOT GIVEN DIRECTLY, WE HAVE TO USE CROSS RATE CONCEPT 67 68 69 70 71 72 73 74 75 76 EUR/AUDEUR/USD X USD/AUD MATHEMATICALLY, USD WILL GET CANCEL WE WILL GET EUR/AUD SO WE MULTIPLY THESE TWO RATES BUT THE PROBLEM IS THAT WE ARE GIVEN AUD/USD & NOT USD/AUD SO WE REARRANGE IT = EUR/USD X USD/AUD = EUR/USD X [1/(AUD/USD)] = 1.199 X (1/0.7980)-1.5025 IT IS JUST BASIC MATHS, WE DID RECEIPROCAL 4 H KE CAPM UTILITY, SH beta bond c M / future INDEX | INTL CAP BUD ING PV, FV, ANNUITY DIR cleanYIELD bond stru WACC RESI ex di 福 130% rences: IN7 13:50 31-01-2019Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text ta copy ▼ ,_a. ars-函Merge & Center, $, % , 弼,8 C Paste B l u. Conditional Format CeInsert Delete Format Formatting, as Table w styles. ▼ ㆆ ▼ Sort &Find & 2 ClearFe Select Edting Format Painter Clipboard Font Alignment Number Styles Cells AR83 AP AQ AR AS AT AU AV AW AX AY AZ BA BC BD BE 78 79 80 81 82 83 84 85 86 87 SECOND ONE: AUD/THB SAME WAY AUD/THB = AUD/USD X USD/THB HERE WE ARE GIVEN BOTH RATES IN SAME FORMAT SO WE SIMPLY MULTIPLY AUD/USD X USD/THB 0.7980 X 32.50 25.9350 89 90 91 92 93 94 95 SING PV, FV, ANNUITYDIR stru WACCRESI ex d 福 130% 4 H KE CAPM UTILITY, SH beta bond c M / future INDEX | INTL CAP BUD cleanYIELD bond erences: 130 13:52 31-01-2019